OpenStudy (anonymous):

Write the expression with positive exponents and simplify as much as possible (For example, instead of 1/9^2, enter 1/81) Simplify: (-2) ^ -1 =

OpenStudy (lucaz):

negative exponents gives you the reciprocal with positiveexponent in the denomitator, so b^-exp=1/b^exp (-2)^-1=1/(-2)